OpenStudy (anonymous):

the perimeter of a rectangle is 32 meters and its area is 48 square meters. Find the length and width of the rectangle.I need help on how u set this up.

jhonyy9 (jhonyy9):

width = 4 lenght = 12 perimeter = 4+4+12+12=32 area =4*12=48

OpenStudy (anonymous):

I just struggle with some of the word problems I have been out of school 12 years and I struggle setting up the problems.I know the formulas but i always get stuck on the relationships between the 2 like combined areas

jhonyy9 (jhonyy9):

how you can make these calcules ? - so you know that the perimeter =2l +2w - the area = w*l - so area=w*l --- 48=w*l --- l=48/w perimeter = 2w +2l 32= 2w + 2(48/w) 32w = 2w^2 +96 2w^2 -32w +96 =0 / 2 w^2 -16w +48 =0 (w-12)(w-4)=0 w-12=0 w=12 w-4=0 w=4 - so if l=48/w than l=48/12 =4 and l=48/4 =12 - so from this result that l=12 and w=4

jhonyy9 (jhonyy9):

- hope so much that is understandably now sure

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

ok im gonna copy that down thankk you for the help i really appreciate it. i have 2 more word problems i hope i can get help with.